Simplify the expression 5/(x-1) – 2/(x+1) over x/(x-1) + 1/(x+1) =A/B
Your answer for A is
Your answer for B is

Answer by Eezzdboy(5)

A=3/x-1 and B=-1+or-sprt2/x^2-1